# Hybrid inflation from supersymmetry breaking

### v1

@@ 1 Introduction
In past works [1, 2, 3], a framework of intimate connection between supersymmetry and inflation was introduced, dubbed inflation by supersymmetry breaking, relating two theoretical proposals motivated, correspondingly, by particle physics and cosmology. The basic idea is to identify the inflaton with the superpartner of the goldstino, charged under a gauged R-symmetry with the corresponding $U(1)$ field becoming massive by absorbing the (complex) inflaton phase which is the R-goldstone boson. The superpotential is forced to be linear in the inflaton superfield $X$ by symmetry, while the Kähler potential can be expanded around its canonical form in powers of $X{\bar{X}}$ with small coefficients representing quantum corrections of the underlying supergravity theory. For a positive sign of the first order correction, neglecting D-term contributions suppressed by a small R-gauge coupling, the scalar potential has a maximum that can realise hilltop inflation with a spectral index of primordial scalar perturbations determined by this correction. Moreover, a second order correction allows for accommodating a nearby minimum with tuneable vacuum energy. The (supersymmetric) Standard Model (SM) can be coupled in a straightforward way with the inflaton part playing also the role of the supersymmetry breaking sector, while the gauge R-symmetry may contain the usual R-parity as a subgroup. The model is very predictive but leads to a superheavy spectrum of superparticles near the inflation scale.
@@
In this work, we generalise the above framework to hybrid inflation [4, 5], by introducing a waterfall direction in the scalar potential that opens up at a critical point at the end of inflation along which the waterfall field falls rapidly to a deep global minimum while the inflaton is displaced slightly. The waterfall field is neutral under the R-symmetry and its superpotential can be adjusted so that the global minimum stays infinitesimally close to a supersymmetric one which consists of a valley around a flat direction for the inflaton field. This allows for a perturbative treatment of the whole dynamics around the origin for the inflaton field and around the supersymmetric minimum for the waterfall with the supersymmetry breaking scale being a free parameter independent of the inflation scale. Moreover, without affecting the inflationary predictions of the model, the presence of the waterfall direction provides a way to end inflation efficiently and tune the vacuum energy of the global minimum at a value infinitesimally close to zero. The cancellation occurs between the negative F-term and positive D-term contributions to the scalar potential. [^1] Previous works on hybrid inflation in the supersymmetric framework can be found for example in [6]. See also [7] and references therein for recent developments.
@@ 2.1.2 Coupling with supersymmetric Standard Model sector
Model 1: $~{}~{}\hat{K}=K+\sum{\zeta}\bar{\zeta},\quad\hat{W}_{1}=X[fw(\phi)+\mu H_{u}H_{d}+y_{u}\bar{u}{\mathcal{Q}}H_{u}-y_{d}\bar{d}{\mathcal{Q}}H_{d}-y_{e}\bar{e}LH_{d}],$ (2.7)
Model 2: $~{}~{}\hat{K}=K+\sum{\zeta}\bar{\zeta},\quad\hat{W}_{2}=X[fw(\phi)+\mu H_{u}H_{d}]+y_{u}\bar{u}{\mathcal{Q}}H_{u}-y_{d}\bar{d}{\mathcal{Q}}H_{d}-y_{e}\bar{e}LH_{d},$ (2.8)
@@ 3 Vacuum structure
Recall that inflation with the subsequent waterfall is supposed to occur for small $\rho$. We can then deal with the potential perturbatively in $\rho$. The strategy of our analysis is that we first find the extrema of the potential at $\rho=0$ as the unperturbed part, and then study how they evolve when $\rho$ is turned on. In the following analysis, we shall use a new parameter ${\delta}$ instead of ${\lambda}$, defined by333 Note that ${\delta}^{2}$ is supposed to be positive, so that the potential at $\rho=0$ is stable in $\phi_{R}$ around $\phi_{R}=0$.
@@ 3.2.1 Analysis in the absense of the D-term potential
where we introduced ${\gamma}^{\prime}={\gamma}+(A\mu/2)$ to absorb $A$444 This can be understood as follows: at the leading order in ${\delta}$, the $X^{2}\bar{X}^{2}$ term in the Kähler potential becomes $(A+2{\gamma}/\mu)X^{2}\bar{X}^{2}$, so $A$ appears in $V_{F}(\rho,\phi_{0}(\rho))$ up to $\rho^{4}$ only through the combination $A+2{\gamma}/\mu$. The stability at the extremum $\phi_{0}(\rho)$ can be seen by expanding the potential there:
@@ 5.2 Particle spectrum
We finish our analysis by presenting as benchmark model an explicit example with low mass spectrum in the multi-TeV region within the reach of next generation of particle colliders. It contains a neutralino mostly electroweak gaugino as lightest supersymmetric particle (LSP) which is a standard Dark Matter candidate and predicts the existence of the universal inflaton sector ($U(1)_{R}$ gauge boson and gaugino, as well as the inflaton) among the low energy spectrum, which is the smoking gun of inflation by supersymmetry breaking. Let us consider Model 2 in Figure 5, where the choice $(\hat{q},\rho_{m}^{2})=(1,1/4)$ satisfies the lower bound (5.11) from the weak gravity conjecture applied to the gravitino. We choose $(x,{\gamma}^{\prime})\simeq(0.1454,0.09318)$, which is the intersection of the $\rho_{m}^{2}=1/4$ and $\rho_{\mathrm{end}}=\rho_{c}$ lines. The constraint on ${\delta}^{2}$ is ${\delta}^{2}\ll 2.4\times 10^{-8}$. For example, we may choose ${\delta}^{2}$ so that the masses of the soft scalars and MSSM gauginos are around $10\sim 100$ TeV. If we choose ${\delta}^{2}\sim 10^{-14}$ and use the values for the coupling constants of the MSSM gauge groups (see for instance [11]),
@@ 6 Conclusion
The requirement of perturbative treatment of both inflation and potential minimum, together with detailed investigation of the parameter space done in Section 4, leads a light spectrum of order $\delta^{2}\ll f$ that contains besides the usual MSSM fields (scalar quarks and leptons and SM gauginos) the universal inflaton sector of inflation by supersymmetry breaking: $U(1)_{R}$ gauge boson and gaugino together with the inflaton. The light mass spectrum has the following approximate pattern: the electroweak gauginos tend to be the lightest as long as shown in the explicit example we presented in Section 5. The next lightest group contains the squarks and sleptons, $U(1)_{R}$ gauge boson and gaugino, and the gravitino, having masses of more or less the same order. Then the inflaton comes next a bit heavier. This is a model independent feature of our framework which connects supersymmetry breaking and inflation in a minimal way. Its embedding in a fundamental theory, such as string theory, remains an open challenging question.

### v2

@@ 1 Introduction
In past works [1, 2, 3], a framework of intimate connection between supersymmetry and inflation was introduced, dubbed inflation by supersymmetry breaking, relating two theoretical proposals motivated, correspondingly, by particle physics and cosmology. The basic idea is to identify the inflaton with the superpartner of the goldstino, [^1] For past related literature, see e.g. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. charged under a gauged R-symmetry with the corresponding $U(1)$ field becoming massive by absorbing the (complex) inflaton phase which is the R-goldstone boson. The superpotential is forced to be linear in the inflaton superfield $X$ by symmetry, while the Kähler potential can be expanded around its canonical form in powers of $X{\bar{X}}$ with small coefficients representing quantum corrections of the underlying supergravity theory. For a positive sign of the first order correction, neglecting D-term contributions suppressed by a small R-gauge coupling, the scalar potential has a maximum that can realise hilltop inflation with a spectral index of primordial scalar perturbations determined by this correction. Moreover, a second order correction allows for accommodating a nearby minimum with tuneable vacuum energy. The (supersymmetric) Standard Model (SM) can be coupled in a straightforward way with the inflaton part playing also the role of the supersymmetry breaking sector, while the gauge R-symmetry may contain the usual R-parity as a subgroup. The model is very predictive but leads to a superheavy spectrum of superparticles near the inflation scale.
@@
In this work we address the problem of scale separation, aiming to decouple the scale of inflation from the one of low energy supersymmetry breaking but staying always within the successful framework of inflation by supersymmetry breaking. To this end, we generalise the above framework to hybrid inflation [16, 17], by introducing a waterfall direction in the scalar potential that opens up at a critical point at the end of inflation along which the waterfall field falls rapidly to a deep global minimum while the inflaton is displaced slightly. The waterfall field is neutral under the R-symmetry and its superpotential can be adjusted so that the global minimum stays infinitesimally close to a supersymmetric one which consists of a valley around a flat direction for the inflaton field. This allows for a perturbative treatment of the whole dynamics around the origin for the inflaton field and around the supersymmetric minimum for the waterfall with the supersymmetry breaking scale being a free parameter independent of the inflation scale. Moreover, without affecting the inflationary predictions of the model, the presence of the waterfall direction provides a way to end inflation efficiently and tune the vacuum energy of the global minimum at a value infinitesimally close to zero. The cancellation occurs between the negative F-term and positive D-term contributions to the scalar potential. [^2] Previous works on hybrid inflation in the supersymmetric framework can be found for example in [18]. See also [19, 20, 21] and references therein for recent developments.
@@ 2.1.2 Coupling with supersymmetric Standard Model sector
Model 1: $~{}~{}\hat{K}=K+\sum{\zeta}\bar{\zeta},\quad\hat{W}_{1}=X[fw(\phi)+\mu_{H}H_{u}H_{d}+y_{u}\bar{u}{\mathcal{Q}}H_{u}-y_{d}\bar{d}{\mathcal{Q}}H_{d}-y_{e}\bar{e}LH_{d}],$ (2.7)
Model 2: $~{}~{}\hat{K}=K+\sum{\zeta}\bar{\zeta},\quad\hat{W}_{2}=X[fw(\phi)+\mu_{H}H_{u}H_{d}]+y_{u}\bar{u}{\mathcal{Q}}H_{u}-y_{d}\bar{d}{\mathcal{Q}}H_{d}-y_{e}\bar{e}LH_{d},$ (2.8)
@@ 3 Vacuum structure
Recall that inflation with the subsequent waterfall is supposed to occur for small $\rho$. We can then deal with the potential perturbatively in $\rho$. The strategy of our analysis is that we first find the extrema of the potential at $\rho=0$ as the unperturbed part, and then study how they evolve when $\rho$ is turned on. In the following analysis, we shall use a new parameter ${\delta}$ instead of ${\lambda}$, defined by444 Note that ${\delta}^{2}$ is supposed to be positive, so that the potential at $\rho=0$ is stable in $\phi_{R}$ around $\phi_{R}=0$.
@@ 3.2.1 Analysis in the absense of the D-term potential
where we introduced ${\gamma}^{\prime}={\gamma}+(A\mu/2)$ to absorb $A$555 This can be understood as follows: at the leading order in ${\delta}$, the $X^{2}\bar{X}^{2}$ term in the Kähler potential becomes $(A+2{\gamma}/\mu)X^{2}\bar{X}^{2}$, so $A$ appears in $V_{F}(\rho,\phi_{0}(\rho))$ up to $\rho^{4}$ only through the combination $A+2{\gamma}/\mu$. The stability at the extremum $\phi_{0}(\rho)$ can be seen by expanding the potential there:
@@ 4 Allowed parameter region
Actually, the number of e-folds ${\Delta}N$ is constrained from below by a value around 34.2 plus a logarithmic contribution depending on reheating temperature [26], $\Delta N\gtrsim 34.2+\ln\left(T_{\text{reh}}/1\,\text{TeV}\right)$, allowing for a wider range of ${\Delta}N$ than our choice above. However, in our setup ${\Delta}N$ appears only through the factor $e^{(1-n_{s}){\Delta}N/2}$ (see eq.(2.22)). Even if we adopt for example ${\Delta}N=40\sim 60$, this factor varies only from 2.2 to 3.3, which does not substantially affect the allowed parameter region we find below.
@@ 5.2 Particle spectrum
We finish our analysis by presenting as benchmark model an explicit example with low mass spectrum in the multi-TeV region within the reach of next generation of particle colliders. It contains a neutralino, mostly electroweak gaugino, as the lightest supersymmetric particle (LSP) which is a standard Dark Matter candidate and predicts the existence of the universal inflaton sector ($U(1)_{R}$ gauge boson and gaugino, as well as the inflaton) among the low energy spectrum, which is the smoking gun of inflation by supersymmetry breaking. Let us consider Model 2 in Figure 5, where the choice $(\hat{q},\rho_{m}^{2})=(1,1/4)$ satisfies the lower bound (5.11) from the weak gravity conjecture applied to the gravitino. We choose $(x,{\gamma}^{\prime})\simeq(0.1454,0.09318)$, which is the intersection of the $\rho_{m}^{2}=1/4$ and $\rho_{\mathrm{end}}=\rho_{c}$ lines. The constraint on ${\delta}^{2}$ is ${\delta}^{2}\ll 2.4\times 10^{-8}$. For example, we may choose ${\delta}^{2}$ so that the masses of the soft scalars and MSSM gauginos are around $10\sim 100$ TeV. If we choose ${\delta}^{2}\sim 10^{-14}$ and use the values for the coupling constants of the MSSM gauge groups (see for instance [25]),
@@
The formula for the soft scalar masses $m_{0}^{2}$ is valid only for the squarks and sleptons, while the masses for the Higgs fields get an additional contribution from the superesymmetric mass term $\mu_{H}H_{u}H_{d}$ in the superpotential (2.8) that introduces an additional mass parameter $\mu_{H}$. It is convenient for the following to introduce $\hat{\mu}_{H}\equiv\mu_{H}/m_{3/2}$ which satisfies:
$\hat{\mu}_{H}=\frac{\mu_{H}}{fw(\phi_{m})}\quad;\quad\phi_{m}\equiv\phi_{0}(\rho_{m})\,,$ (5.15)
where $\phi_{m}$ is the vacuum expectation value of the waterfall field at the supersymmetry breaking minimum of the scalar potential (3.24). The Higgs masses and mixing are then given by:
$m_{H_{u}}\simeq m_{H_{d}}\simeq m_{3/2}(1+\hat{\mu}_{H}^{2})^{1/2}\quad;\quad B\mu_{H}=\hat{B}\hat{\mu}_{H}m_{3/2}^{2}\quad;\quad\hat{B}\simeq 10\,,$ (5.16)
where we kept only the dominant contribution, neglecting higher order corrections in the Kähler potential and corrections from $U(1)_{R}$ gauge interactions. Obviously, these contributions are important when $\mu_{H}$ is large, i.e. $\hat{\mu}_{H}$ of order unity, which may be required in some models for the avoidance of the cosmological moduli problem (see e.g. [28]).
@@ 6 Conclusion
The requirement of perturbative treatment of both inflation and potential minimum, together with detailed investigation of the parameter space done in Section 4, leads a light spectrum of order $\delta^{2}\ll f$ that contains besides the usual MSSM fields (scalar quarks and leptons and SM gauginos) the universal inflaton sector of inflation by supersymmetry breaking: $U(1)_{R}$ gauge boson and gaugino together with the inflaton. The light mass spectrum has the following approximate pattern: the electroweak gauginos tend to be the lightest as long as shown in the explicit example we presented in Section 5. The next lightest group contains the squarks and sleptons, $U(1)_{R}$ gauge boson and gaugino, and the gravitino, having masses of more or less the same order. Then the inflaton comes next a bit heavier. This is a model-independent feature of our framework which connects supersymmetry breaking and inflation in a minimal way. Its embedding in a fundamental theory, such as string theory, remains an open challenging question.